An LC circuit is built with an inductor L = 9.40 mH and a capacitor C = 17.9 pF. If the capacitor voltage has its maximum value of V = 4.49 V at t = 0 s, what is the inductor current if the capacitor is fully discharged?
Answer #1

Here ,

let the inductor current is I

for the energy stored in the inductor to be maximum

0.50 * CV^2 = 0.50 L * I^2

17.9 *10^-12 * 4.49^2 = 9.4 * 10^-3 * I^2

solving for I

I = 1.96 *10^-4 A

the current at that time in the inductor is 1.96 *10-4 A